OpenStudy (anonymous):

For what values of k is f(x) continuous? 3x^2-11x-4 x<=4 kx^2-2x-1 x>4

OpenStudy (bahrom7893):

when x = 4, f(x) must be 0

OpenStudy (bahrom7893):

k*4^2-2*4-1=0

OpenStudy (bahrom7893):

there's an alternative method...

OpenStudy (bahrom7893):

which is kinda what i did, but i skipped steps

OpenStudy (bahrom7893):

3x^2-11x-4, when x<=4.. plug in 4 into that equation: 3*4^2-11*4-4=3*16-44-4=48-44-4=0

OpenStudy (bahrom7893):

Thus the second equation must also be 0 when x=4

OpenStudy (bahrom7893):

 lol btw that's not what the functions look like, but you have to make them touch each other